# Ina K. Schieferdecker Technische Universität Berlin, Einsteinufer 25, 10587 Berlin, Germany

### v2

@@ 2 Models in Software Engineering
    In addition to explicit model development and top-down model use, models can also be extracted from software executions, from software monitorings or from code artefacts with the objective of avoiding the costs associated with model development on the one hand and leveraging the benefits of MDSE on the other hand [7].
@@
    The combination of model development of model artefacts with model extraction from code artefacts has resulted in an hybrid top-down/bottom-up MDSE approaches [76] benefiting from both directions of MDSE as shown in [83, 85] or [37].
@@
    Furthermore, as the abstractions inherent in programming became more intricate, low-level information and structure models have become indispensable to conventional programming. This phenomenon is exemplified by the utilisation of data structures that are frequently defined in JSON, XML schemata, or SQL table specifications, to name but a few.
These practical outcomes of model-driven, model-based and/or model-oriented software developments established a foundation for the utilisation of AI in software engineering, as discussed in Section 5. Before that, however, section 3 will take a closer look at the current state of big code on open source platforms, and section 4 will take a closer look at the state of applying AI to big code.
3 Big Code in Software Engineering
The advent of software coding platforms can be traced back to the 1990s, although it was not until the early 2000s that they truly gained prominence, with the launch of GitHub in April 2008 representing a significant milestone in this regard. At the present time, GitHub is the most widely utilised software coding platform. As stated by [15], GitHub hosts over 100 million projects and 40 million users. It became not only a significant platform for software engineering collaboration, but also a prominent reference for open-source software mining [45]. The study also demonstrated that a considerable number of GitHub repositories are not directly related to software development. This is because GitHub is not solely utilised for coding collaborations; it is also employed for collaborations on websites, editing books or other publications, and is even used as a storage platform. A later manual analysis by the same authors [46] revealed that over a third of the repositories on GitHub were not software development repositories, which degraded now to roughly 20 percent according to [4] (see figure 2 upper left).
Nevertheless, the software engineering community can further improve its understanding of software engineering processes, methods and tools by gaining insights through the mining of coding platforms using tools such as Kibble [4]. In a Feb. 2025 Kibble snapshot analysing lines of code from all GitHub code repositories, about 70% is actual code, while the other 30% is comments and empty lines. Java is the most common coding language, followed by XML and JSON for data and data structures (see figure 2 below left). A GitHub search counting files in GitHub repositories shows the ratio of the number of files for coding files to model files for software designs, interface models and data models (see figure 2 top right). The split between design models, interface models and data models is shown in figure 2 below right, which shows that about 1% of the files on GitHub contain models or partial model information.
All of these figures are only rough indicators: Lines of code and number of files are crude measures of software. File extensions do not accurately indicate file types. Coding and modelling tools are numerous and sometimes have proprietary file extensions. Overall, a GitHub search for more specific file types such as UML works better than for more general file types such as XML or JSON, but can still lead to false positives.
Nevertheless, the sheer numbers show that modelling is being used in software development. A more detailed approach to modelling on GitHub, including image analysis, is given in [72], but this is not applicable to GitHub as a whole due to API limitations. Despite this, the numbers show that modelling is being actively used. In fact, modelling is mainly used for data/interface structures, due to the simple fact that data and interfaces do not work without at least a well-defined structure. In contrast, design and behaviour models for complete software (packages) are only used for a small fraction.
This is very much in line with the analysis of [34] over a decade ago: Code generators may produce a large amount of code, but not all of it is necessarily useful or would have been written manually. The effort required to develop, maintain and adapt the generators reduces the productivity gains from modelling. When manual code changes are required, they make round-trip engineering difficult. Generated code must be checked by developers to ensure correctness, limiting the benefits of automation. Poor software design models can lead to inefficient development. As a result, developers often prefer generic solutions with hand-written, reusable code to customised, generated solutions. As programming languages and frameworks become more expressive, the need for model-driven code generation diminishes. In this sense, GenAI further empowers the software developer. But does this mean that MDSE will disappear, or will the ability to (re)generate models from code and/or keep them in sync with software requirements evolution and manual code changes result in a boost for MDSE (see section 5)?
Back to big code: Due to the large number of software repositories with roughly 500 MB code each, the concept of "Big Code", analogous to the notion of "Big Data", was introduced in reference to the vast quantities of software code that have accumulated over time [58, 61, 84]. This code can be reused, extended and analysed to gain deeper insights into software engineering. It can be used to train ML models and enabled the application of AI for SE in large scale [53].
4 Artificial Intelligence in Software Engineering
The use cases for AI in SE are numerous and cover a wide spectrum as discussed already back in 1988 [11] and since then [2]. For example, [31] proposes the AI in SE Application Levels (AI-SEAL) taxonomy, which differentiates between the point of applying AI to the software engineering process, the software product or at runtime, the levels of automation, and the types of AI along the five tribes differentiation by [27]. It is noteworthy that the majority of the papers analysed in [31] employ AI in the software engineering process. However, despite the numerous facets of the software engineering process (Figure 1), this study does not provide further elaboration. Nevertheless, other publications offer more detailed discussions of AI in SE, including [10] or [62]. Notwithstanding the aforementioned considerations, a taxonomy for the role of AI in software engineering has yet to emerge.
In light of the recent advancements in machine learning (ML) that have made significant breakthroughs possible, this paper focuses mainly on the latest developments in the application of ML to SE. For ML to be effective, it is essential to utilise the appropriate structures inherent to SE artefacts, which has been a topic of considerable debate: [3] presents an overview of the various ways in which source code can be represented, including representational models of tokens, token contexts, program dependency graphs, API calls, abstract syntax trees, object usage, and others. These representational models are used for AI assistance in SE, including the creation of recommender systems, the inference of coding conventions, the detection of anomalies and defects, the analysis of code, the rewriting and translation of code, the conversion of code to text for the purposes of documentation and information retrieval, and the synthesis and general generation of code from text. [47] adds further applications of AI assistance such as code completion, API migration and code repair.
Furthermore, [3] presents "[t]he naturalness hypothesis[:] Software is a form of human communication; software corpora have similar statistical properties to natural language corpora; and these properties can be exploited to build better software engineering tools.". Given the intrinsic formats and formal characteristics of coding and modelling languages employed in SE, the second aspect of the naturalness hypothesis can be considered relatively straightforward. Nevertheless, the initial proposition reinforces the necessity for SE artefacts that are readily comprehensible. This assertion was previously made by [33] in a different form: "[A]ny fool can write code that a computer can understand, good programmers write code that humans can understand.". Consequently, the application of AI in SE is not merely concerned with code generation or code repair; it also encompasses the enhancement of code through techniques such as refactorings, with the objective of optimising readability and maintainability.
Figure 3: AI-assisted Coding
Figure 4: AI4SE - The AI4SE Taxonomy of AI-assisted Software Engineering
The purpose dimension has the three facets of (1) understanding to gain insight, (2) generating to derive (parts of), and (3) improving to advance SE artefacts and/or SE processing. An approach may address multiple purposes. If it is hybrid in this sense, all its purposes are made explicit, e.g. to support understanding only is denoted by ’1’, or to support both understanding and generation is denoted by ’1, 2’.
The AI4SE application targets consist of the main SE activities as described by SWEBOK [73] and DevOps [51]: (1) Software development includes requirements engineering, architecture and design, coding, testing, and maintenance. (2) Software operations includes deployment, execution, and monitoring. Together with the (3) SE process including management and adjustment, the three form the target of application dimension in the taxonomy. Each of the target options can be performed with the support of models, based on models or even driven by models, or without them. For the sake of comprehension of the taxonomy, the model option is optional for each application target: If models are used, i.e. the application target is MDSE, it is marked with "(m)". Thus, the application target "1.m" refers to software development with models, and "1" to one without. Or, in a more fine-grained analysis, "1.d.m" would be used for model-based testing and "1.d" for testing without models. An approach may address several application targets. If it is hybrid in this sense, all of its goals can be made explicit, e.g. support for coding only is denoted by "1.c", support for both coding and testing is denoted by "1.c, 1.d", or support for coding and model-based testing is denoted by "1.c, 1.d.m". For simplicity, "h" can also be used to indicate a hybrid approach of several targets.
The types of AI are divided into (1) symbolic AI, i.e. knowledge representation, rule-based systems, and logical reasoning, (2) statistical AI, i.e. probabilistic modelling and inference, Bayesian networks, Markov models, Gaussian processes, etc., (3) classical machine learning, i.e. supervised, unsupervised, and reinforcement learning, feature engineering and selection, etc., (4) deep learning, i.e. neural networks, transformers, transfer learning, self-supervised learning, attention mechanisms, etc., (5) generative AI (GenAI), i.e. large language models (LLM), generative adversarial networks, diffusion models, variational autoencoders, autoregressive models, contrastive learning, reinforcement learning from human feedback, etc., (6) agentic AI, i.e. multi-agent systems, planning and decision making, hierarchical reinforcement learning, etc., and (7) general purpose AI, i.e. meta-learning and world models. Today, all types (1) to (6) are used in software engineering, but not (7) due to its non-existence. An approach can also use a combination of AI types, i.e. it can be hybrid. This can be made explicit, e.g. by denoting "1, 2" for a hybrid approach that uses both symbolic and classical machine learning. For simplicity, "h" can just be used to denote a hybrid approach that uses more than one type of AI.
We define the levels of autonomy/automation similarly to the levels of autonomous driving [9] as five levels from no support to recommendations to full automation: (1) none, (2) AI-assisted options, where the developer is in full control and receives recommendations to choose from, (3) AI-assisted selection, where the AI pre-selects options, (4) AI-based partial automation, where the AI selects options in simple, standard cases, and (5) AI-based full automation, where the AI operates without the developer. So far, level 3 is the most common, level 4 is on the rise and level 5 is rather unrealistic for complex, industrial-scale software. Higher levels of autonomy typically include the capabilities of lower levels of autonomy; when presenting the level of autonomy of an approach, the highest level of autonomy is given.
Paper Summary Purpose Target Type Level
[55] T-BERT generates trace links between source code and natural language artefacts. It represents a significant advancement in automated software traceability, offering improved accuracy, efficiency, and practicality. 2, 3 1 4 4
[49] NLP approach for improving requirements engineering in the automotive industry: It identifies flaws, such as ambiguous words, incomplete process descriptions, or incorrect quantifiers. 3 1.a 1 3
[63] CBRank method for prioritising software requirements: It uses stakeholder input and requirement attributes to generate an optimal ranking and enables adaptive elicitation by refining the prioritisation based on real-time feedback. 3 1.a, 2.a, 3.1 1, 3 3
[12] ADeX is a tool designed to automate the curation of architectural design decisions in software development. It helps automate the management of architectural decisions, reducing documentation effort while improving decision accuracy. 1, 3 1.b, 2.a 1, 3 4
[14] Explores the early experiences of developers using the AI-augmented coding assistant GitHub Copilot. It may may expand to debugging, code review, and software maintenance, requiring better trust and provenance tracking. 1, 2, 3 1.c 5 4
[78] IntelliCode Compose is a code completion tool that uses a generative transformer model to predict sequences of code tokens, including entire lines of code. It has been trained on a massive dataset of 1.2 billion lines of code across multiple programming languages. 2, 3 1.c 2, 4, 5 4
[6] Getafix is an automated bug-fixing tool that learns from past, human-written fixes to suggest human-like fixes for static analysis warnings. 2, 3 1.c, 1.d 2, 3 3
[38] GraphCodeBERT improves code understanding by incorporating dataflow graphs into the BERT transformer architecture. Data flow graphs represent dependencies between code variables and improve semantic understanding over treating code as a simple sequence of tokens. 1, 3 1.c, 1.d, 1.e 2, 4, 5 4
[40] DeepCodeReviewer is an automatic code analysis system that learns from historical peer reviews to identify and suggest relevant reviews for code snippets. 1, 3 1.c, 1.d, 3.a 2, 3, 4 4
[28] Introduces DeepDebug or automated bug detection and fixing in Java: It generates non-deletion fixes that modify instead of just deleting code 3 1.c, 1.e 4, 5 4
[81] Approach improves code review automation with DL by working with pre-trained models on raw source code and English text. 3 1.c, 1.d 5 3
[79] The book discusses the application of AI beside others for infrastructure monitoring and anomaly detection, predictive maintenance, and build pipeline optimization. 1, 3 1.c, 1.d, 2.a, 3.b 2, 3 4
[16] Presents RepairAgent for automated program repair that allows to autonomously plan and execute actions, such as gathering information, searching for code snippets, and validating fixes, by interacting with a set of specifically designed tools. 3 1.c, 2.a 5, 6 4
[52] Approach links different types of testing information to help automate the generation of equivalence classes for testing. The linking supports tasks such as test reduction, prioritisation and selection. 1, 2, 3 1.d 2, 3 3
[87] Presents a hybrid approach to test case prioritization for regression testing aiming to minimise testing costs. Aims to enhance software testing efficiency by optimising test case execution order. 3 1.d, 1.e 1, 2, 3 3
[75] Introduces Retecs for automatic test case prioritisation and selection in continuous integration. It uses test case meta-data to adapt to changes in the environment and prioritise error-prone test cases. 3 1.d, 2.a 3, 4 4
[21] Explores machine learning-based ensemble methods to predict software quality more accurately than previous approaches 1, 3 1.d, 3.a 3 4
This novel AI4SE taxonomy [^2] If you want to work on it further, please contact the author for the ontology file which further refines the targets of applications and the AI types. is shown in Figure 4. The fidelity of this new taxonomy AI4SE is verified by reviewing and classifying highly cited publications, but also recent publications to demonstrate the applicability of this taxonomy. The classification of the papers is given in table 1. Additional details on the papers that target mainly the software development are given here:
@@ 4 Artificial Intelligence in Software Engineering
For requirements, [49] presents an AI-based automation approach to requirements engineering that begins by converting natural language into an Eclipse Modelling Framework (EMF) model. It then applies linguistic rules to identify errors, such as ambiguities or incorrect quantifiers, and provides suggestions for requirements analysts to make final decisions. This approach supports the entire requirements elicitation and change process. Another requirements improvement is given in [63] presenting the prioritisation of requirements by combining the preferences of project stakeholders with approximations of the order of requirements computed by ML techniques.
@@
For architecture & design, [12] describes the automated curation of design decisions to support architectural decision-making. It helps software architects by organizing and recommending design decisions based on previous cases and contextual information of the current project. By leveraging existing design knowledge, the approach analyses historical data and design choices to improve the quality and consistency of architectural designs.
@@
Overall, the majority of the presented selection of recent research focuses on improving the software engineering process, either by improving coding, requirements engineering, testing, or the overall processes. A smaller number of approaches aim to improve the understanding of software engineering artefacts. There are also methods that focus on generating code or code-related content. A mixture of AI types are used, including symbolic, statistical, ML, DL and GenAI, as well as hybrid approaches. Many of the AI approaches offer AI-assisted selection, i.e. they provide suggestions, recommendations and rankings that can help developers, but require a developer to make the final decision on what to do. AI-based partial automation is also a common level, where tools automate part of the process, but may require a human to review or further integrate the results. The facets of the AI4SE taxonomy are adequately addressed, although not all options are covered. Nevertheless, it is only a matter of time before research results are presented in primary studies for the open combinations of the facets, as long as they provide advances beyond the state of the art.
@@ 5 AI-Assisted Model-Driven Software Engineering
The substantial increase in the utilisation of AI in SE is in parallel accompanied by a notable rise in the application of AI to MDSE. This phenomenon can also be explained by our novel modelling hypothesis, which is an extension of the naturalness hypothesis [3] to modelling artefacts in software engineering:
The modelling hypothesis
Software is (also) a formalized communication between humans and computers; model corpora, software corpora, and natural language corpora have similar statistical properties; the properties of model corpora can be employed to develop more efficacious software engineering tools.
Indeed, the advent of big code and AI has opened up new avenues for integrating AI into model-driven approaches, as posited by the novel modelling hypothesis. As demonstrated by [42], the intrinsic formal graph structures of SE models can also be leveraged to facilitate the preparation of models for gaining AI advantages in MDSE. This is analogous to the considerations in [64] on the necessity to elevate abstraction levels in language models to enhance the efficacy of AI-assisted automated coding.
@@
[50] Discusses NLP-driven model generation for more efficient model construction. Presents "purposive" meta-models that guide the interactions between domain experts and GenAI to generate effective prompts for model construction. 2 1.a.m, 1.b.m 1, 5 4
[22] Discusses the NLP-driven domain modelling by evaluating different LLMs and various prompt engineering techniques. 2 1.a.m, 1.b.m 5 4
[80] Details the creation and validation of a high quality dataset of UML diagrams. Provides a valuable research resource in the form of the UML dataset and presents a new, improved approach to UML diagram classification. 3 1.a.m, 1.b.m, 1.c,m, 1.d.m 4 3
[72] Presents a deep learning approach for automatically classifying images of UML diagrams. 1 1.b.m 4 3
[8] Presents an approach to automatically generate model transformations from examples. The approach is demonstrated through experimental evaluation on seven diverse transformation problems. 1, 2 1.b.m 1, 2, 3 4
[57] Explores the automated classification of UML class diagrams into forward-engineered and reverse-engineered diagrams. It provides a dataset with ground truth for the classification of class diagrams. 1, 3 1.b.m 3 3
[1] Presents an experience report detailing the use of LLMs to analyse and transform software variants represented in various formalisms (Java, UML, GraphML, statecharts). 1, 2 1.b.m, 1.c.m 5, 6 3
[29] Presents a software architecture for the AI-augmented automation of DevOps pipelines to address both architectural and technical challenges associated with the development and operation of complex systems. 1 1.b.m, 1.e.m 4 3
[5] Proposes a approach for statistically analysing large collections of models within the field of MDSE. The approach demonstrates its effectiveness in domain analysis and repository management, and shows the potential for applications such as model clone detection. 1 1.b.m, 1.e.m 2, 3 2
[54] Introduces FlowGen, a code generation framework that leverages multiple LLM agents to emulate different software process models (Waterfall, TDD, and Scrum). On standard benchmarks, the Scrum-based FlowGenScrum consistently outperforms other models and baselines. 2 1.c.m, 1.d.m, 1.e.m, 3.1.m, 3.2.m 5, 6 4
@@
As shown in Table 2, the growing trend in the adoption of AI for MDSE is reflected by various new research results:
@@
Foundational work on ground truths for developing and benchmarking AI applications in MDSE is discussed in [80, 72] and [57]. [80] presents the creation and validation of a new, high-quality ground truth dataset of UML diagrams. The curated dataset of 2626 unique and accurately labelled images across six UML diagram types goes beyond the Lindholmen dataset of modelling artefacts [66, 67]. The new ground truth dataset was also presented for an improved approach to UML diagram classification. Also [72] presents a new dataset of images, manually labelled into four UML diagram types (class, activity, sequence, use case) and a non-UML category, which is used to develop new DL approaches with higher precision to automate UML diagram classification. Another ground truth dataset is described in [57]: A manually labelled subset of the Lindholmen dataset is used to explore the automated classification of UML class diagrams into forward-engineered (FwCD) and reverse-engineered (RECD) diagrams using supervised machine learning. The methodology in [5] uses techniques from information retrieval, NLP, and ML to enable an efficient comparison, clustering and clone detection of model artefacts.
[50] presents a methodology for AI-augmented MDSE as a continuous cycle of AI-generated model proposals by LLMs that are then reviewed and refined by humans. It seeks to bridge the gap between the natural language used by domain experts and the formal languages used in MDSE.
[22] investigates the feasibility of using LLMs to automate domain modelling in SE. It applies various prompt engineering techniques to a dataset of diverse domain modelling examples and finds that while LLMs show promising domain understanding and achieve relatively high precision in generating classes and attributes, their recall remains low, particularly for relationships. It concludes that full automation is currently impractical.
@@
[8] presents a novel method for automatically generating model transformations from examples. It employs a three-step process: (1) analysing example model pairs to identify groups of similar transformations, (2) using genetic programming to learn basic rules for each group, and (3) refining these rules using simulated annealing to handle complex dependencies and value derivations. The method presented improves the efficiency and accuracy of learning complex model transformations.
[1] explores the use of LLMs to aid in the complex process of re-engineering existing software variants into a software product line (SPL) by identifying features to represent variability and commonalities. It describes the methodology used to prompt the LLMs, the results obtained, and the limitations encountered. There is a potential in assisting domain analysts and developers in transforming software variants into SPLs, but LLMs are limited in managing large variants or complex structures.
[29] presents a novel software architecture designed to address the increasing complexity of modern system development within a DevOps framework. The architecture integrates MDSE and AI techniques to enhance the entire software lifecycle, from requirements engineering to monitoring and maintenance. The architecture is evaluated through multiple industrial case studies.
@@
[54] uses agentic AI where each agent embodies a specific development role such as requirement engineer, developer, or tester and collaborates via chain-of-thought prompting. The set of LLM agents emulate different software process models and are evaluated against four benchmarks (HumanEval, HumanEval-ET, MBPP, and MBPP-ET). The impact of individual development activities on code quality is also investigated, revealing the critical role of testing and the beneficial effects of design and code review.
@@
So when AI and MDSE are combined, it becomes clear that there are significant opportunities:

### v3

@@ 4 Artificial Intelligence in Software Engineering
      To determine research an AI for SE
@@
      To test the validity of the ai4se taxonomy with a research selection.
@@
      To determine a classification of the selected research.
  • Research questions:
    • RQ1: What is the current status of research on AI for SE and for MDSE?
    • RQ2: Are the dimensions of the ai4se taxonomy suitable for classifying current AI for SE research results?
    • RQ3: Are all dimensions and facets of the taxonomy represented in research work?
    • RQ4: What are typical purposes addressed, targets tackled, types used, and levels of automation with AI reached?
  • Database: An automated search on dblp, ACM DL, and IEEE Xplore of SE and AI related research papers published between 2020 and 2025. Snowballing extension of the search results with Research Rabbit, Citation Gecko, and Google Scholar. Consolidation of the search results with Zotero.
  • Inclusion criteria:
    • Original research papers in English.
    • Online available.
    • Research on AI for SE.
  • Exclusion criteria:
    • Research on SE for AI.
  • Selection process:
    1. 1.
      Title and abstract screening for the pre-selection of unique research candidates by use of the concept map resulting from the ai4se ontology including synonyms.
    2. 2.
      Full text review and assessment of the research contributions for the final selection of unique research.
    3. 3.
      Tools: Python for text analysis and post-processing of finally selected research, supported by MS Visual Studio, Google AI Studio, and LibreOffice.
  • Synthesis process:
    • Review of the finally selected research and inclusion in the ai4se ontology.
A total of 229 studies on AI for SE were identified. Of these, 32 were primarily SLRs, 46 papers described (elements of) research roadmaps for (aspects of) AI for SE, and 10 papers presented taxonomies for (aspects of) AI for SE. Twelve SLRs, eight roadmaps and seven taxonomies are summarised and compared in Tables 3, 4 and 5 respectively, due to their relation to the ai4se taxonomy. A further 82 papers present original methods, techniques, and tools for AI for SE excluding AI for MDSE. Of these, 17 were finally selected and classified as shown in Table 1. For research on AI for MDSE, see Section 5 for details of the 16 research papers selected from a total of 59 research items on AI for MDSE. Please note that the SLR results are available online, and that the ai4se taxonomy will be extended in the future to include further original research papers.
@@
The selected original research on AI for SE in general is presented in Table 1. Additional details are given here: For requirements, [67] presents an AI-based automation approach to requirements engineering that begins by converting natural language into an Eclipse Modelling Framework (EMF) model. It then applies linguistic rules to identify errors, such as ambiguities or incorrect quantifiers, and provides suggestions for requirements analysts to make final decisions. This approach supports the entire requirements elicitation and change process. Another requirements improvement is given in [88] presenting the prioritisation of requirements by combining the preferences of project stakeholders with approximations of the order of requirements computed by ML techniques.
@@
For architecture and design, [20] describes the automated curation of design decisions to support architectural decision-making. It helps software architects by organizing and recommending design decisions based on previous cases and contextual information of the current project. By leveraging existing design knowledge, the approach analyses historical data and design choices to improve the quality and consistency of architectural designs.
@@
Overall, the majority of the presented selection of recent research focuses on improving the software engineering process, either by improving coding, requirements engineering, testing, or the overall processes. A smaller number of approaches aim to improve the understanding of software engineering artefacts. There are also methods that focus on generating code or code-related content. A mixture of AI types are used, including symbolic, statistical, ML, DL and GenAI, as well as hybrid approaches. Many of the AI approaches offer AI-assisted selection, i.e., they provide suggestions, recommendations and rankings that can help developers, but require a developer to make the final decision on what to do. AI-based partial automation is also a common level, where tools automate part of the process, but may require a human to review or further integrate the results. On the whole, the dimensions of the ai4se taxonomy are adequately covered, although not every possible combination has yet been addressed — it may even be impossible to do so. For example, the prospect of full automation driven by AI using only symbolic AI is unlikely, since the nuances of SE can only be effectively managed with more advanced non-symbolic AI. However, as AI for SE is a very active area of research, further research results will emerge that make use of the open combinations of the taxonomy dimensions. To accommodate these results, the taxonomy is available online [96] and will be continuously updated.
@@ 5 AI-Assisted Model-Driven Software Engineering
The substantial increase in the utilisation of AI for SE, as described in the previous section, is in parallel with a notable rise in the application of AI to MDSE in the following areas: model creation and elicitation; model completion, refinement and repair; intelligent model transformations; model validation and verification; reverse engineering and model discovery; and model-based testing. Using AI in MDSE enables the management of unprecedented complexity, improving the quality and consistency of software projects. These developments can also be explained by the novel ’model naturalness hypothesis’, which extends the naturalness hypothesis [8] to modelling artefacts in software engineering:
The model naturalness hypothesis
Software is a form of human communication as well as a form of human-computer communication; model corpora, software corpora, and natural language corpora have similar statistical properties; the properties of model corpora can be employed to develop more efficacious software engineering tools.
Readers familiar with software modelling techniques will not be surprised by this hypothesis. Like software code, software models are expressed in (semi-)formally defined languages with precise, often graphical, syntax and semantics. In this respect, software models are similar to software code. However, they are often only used in the initial phases of software design to sketch the desired solution. Consequently, they are used to visualise code and are not stored in a machine-readable format for reuse and further processing. In contrast, however, software models can facilitate human-human and human-computer communication if they are treated as machine-processable, high-level software artefacts rather than images (see also [12]). In this case, models can form corpora that can be analysed automatically and used to train AI. The model naturalness hypothesis has therefore been made explicit. Indeed, the advent of big code and AI has opened up new avenues for integrating AI into model-driven approaches. As demonstrated by [53], the intrinsic formal graph structures of SE models can also be leveraged to facilitate the preparation of models for gaining AI advantages in MDSE. This is analogous to the considerations in [89] on the necessity to elevate abstraction levels in language models to enhance the efficacy of AI-assisted automated coding.
@@
[12] This paper presents quantitative evidence from academia and industry regarding the increased use of models. It suggests using models as data and applying machine learning techniques to improve our understanding of them. P1 S1.m T2.b L0
[64] Discusses NLP-driven model generation for more efficient model construction. Presents "purposive" meta-models that guide the interactions between domain experts and GenAI to generate effective prompts for model construction. P2 S1.a.m, S1.b.m T1, T3 L3
[30] Discusses the NLP-driven domain modelling by evaluating different LLMs and various prompt engineering techniques. P2 S1.a.m, S1.b.m T3 L3
[3] This paper introduces SimIMA, an intelligent modelling assistant for Simulink. It uses association rule mining and model clone detection to support the modeller in developing models. P3 S1.b.m T2.b L1
[122] This study suggests DDA, an approach to domain-knowledge-integrated, designable data augmentation with transfer learning, which improves the efficiency of deep learning-based metamodelling for complex systems design, such as bumper design for vehicle crash safety. P2 S1.b.m T2.c L1
[93] This paper proposes an approach to bot-modeller interaction. It combines an incremental learning strategy with the discovery of alternative configurations, improving the accuracy of the bot’s suggestions through the analysis of domain modelling decisions over time. P2, P3 S1.b.m T2.b L1
[63] This paper classifies Ecore metamodels by using their structural and textual information to improve automated model classification for model management. P1 S1.b.m T2.a L0
[54] This study examines an approach to improving domain modelling by incorporating micro-learning into the modelling process. P3 S1.b.m T2.a L0
[15] Presents an approach to automatically generate model transformations from examples. The approach is demonstrated through experimental evaluation on seven diverse transformation problems. P1, P2 S1.b.m T1, T2.a, T2.b L3
[76] Explores the automated classification of UML class diagrams into forward-engineered and reverse-engineered diagrams. It provides a dataset with ground truth for the classification of class diagrams. P1, P3 S1.b.m T2.b L2
[2] Presents an experience report detailing the use of LLMs to analyse and transform software variants represented in various formalisms (Java, UML, Graph Markup Language (GraphML), statecharts). P1, P2 S1.b.m, S1.c.m T3, T4 L2
[40] Presents a software architecture for the AI-augmented automation of DevOps pipelines to address both architectural and technical challenges associated with the development and operation of complex systems. P1 S1.b.m, S1.e.m T2.c L2
[11] Proposes an approach for statistically analysing large collections of models within the field of MDSE. The approach demonstrates its effectiveness in domain analysis and repository management, and shows the potential for applications such as model clone detection. P1 S1.b.m, S1.e.m T2.a, T2.b L1
[121] This work proposes Code Generation By Example (CGBE), a code generation approach that has been learned from examples for UML to Java code generation. P2 S1.c.m T1 L1
[70] Introduces FlowGen, a code generation framework that leverages multiple LLM agents to emulate different software process models (Waterfall, Test-Driven development (TDD), and Scrum). On standard benchmarks, the Scrum-based FlowGenScrum consistently outperforms other models and baselines. P2 S1.c.m, S1.d.m, S1.e.m, S3.1.m, S3.2.m T3, T4 L3
[107] This study presents an approach to deriving feature models in UML from historical source code, which can then be used as a basis for generating and prioritising test cases. P2 S1.d.m T2.b L2
@@
Foundational work on ground truths for developing and benchmarking AI applications in MDSE is discussed in [110, 99] and [76]. [110] presents the creation and validation of a new, high-quality ground truth dataset of UML diagrams. The curated dataset of 2626 unique and accurately labelled images across six UML diagram types goes beyond the Lindholmen dataset of modelling artefacts [91, 92]. The new ground truth dataset was also presented for an improved approach to UML diagram classification. Also [99] presents a new dataset of images, manually labelled into four UML diagram types (class, activity, sequence, use case) and a non-UML category, which is used to develop new DL approaches with higher precision to automate UML diagram classification. Another ground truth dataset is described in [76].
The selected original research on AI for MDSE is categorized in Table 2. Further details can be found here: The majority of the original work on AI for MDSE is focuses on supporting the design of software: [12] was one of the first studies to suggest treating models as data for training LLMs, enabling the application of ML techniques to MDSE.
Another significant area of research concerns the use of AI to improve domain modelling. [64] presents a methodology for AI-augmented MDSE as a continuous cycle of AI-generated model proposals by LLMs, which are then reviewed and refined by humans. This approach aims to bridge the gap between the natural language used by domain experts and the formal languages employed in MDSE. [31] investigates the feasibility of using LLMs to automate domain modelling in SE. It applies various prompt engineering techniques to a dataset of diverse domain modelling examples and finds that while LLMs show promising domain understanding and achieve relatively high precision in generating classes and attributes, their recall remains low, particularly for relationships. It concludes that full automation is currently impractical. In [122], the presented method, DDA, was shown to enable efficient performance prediction and design variable estimation by generating virtual responses from a small set of actual responses and integrating domain knowledge. When applied to vehicle bumper design, the method produced virtual responses that were approximately 95% similar to real data, outperforming existing metamodels.
[93] addresses the question of how to improve the modeller–bot interaction. The bot presented in the study suggests alternative model configurations based on a novel algorithm and updates the domain model based on modeller input. This addresses the limitations of existing automated methods. Evaluations demonstrate high accuracy and rapid performance. [54] also examines the interaction between the modeller and the AI tool and propose a multistrategy learning approach for the gradual, goal-driven refinement of incomplete domain models provided by human experts. The system learns general rules from specific examples and refines or extends concepts in order to handle exceptions caused by model incompleteness. An intelligent modelling assistant for Simulink is presented in [3]. It offers two forms of data-driven guidance: SimGESTION provides real-time edit suggestions using ensemble learning, while SimXAMPLE recommends related models through clone detection.
Model management is central to MDSE and is therefore an active area of research. The methodology in [11] uses techniques from information retrieval, NLP, and ML to enable the efficient comparison, clustering and clone detection of model artefacts. The automated classification of UML class diagrams into forward-engineered and reverse-engineered diagrams presented in [76] provides another perspective on model management that is crucial for understanding the evolution of software projects. Different to that, [63] focuses on Ecore metamodels and presents results showing that incorporating structural information significantly improves the accuracy with which models are classified, which in turn helps to improve the comparison, clustering, and clone detection of model artefacts.
@@
Model transformation has also been studied: [15] presents a novel method for automatically generating model transformations from examples. It employs a three-step process: (1) analysing example model pairs to identify groups of similar transformations, (2) using genetic programming to learn basic rules for each group, and (3) refining these rules using simulated annealing to handle complex dependencies and value derivations. The method presented improves the efficiency and accuracy of learning complex model transformations.
@@
Code generation supported by models and AI has also been a key research field: [121] aims to improve the synthesis of reusable and adaptable code generators in MDSE by reducing the manual effort typically required. The proposed approach, Code Generation By Example (CGBE), uses symbolic machine learning and tree-to-tree mappings to learn code generation from UML to Java. [70] uses agentic AI where each agent embodies a specific development role such as requirement engineer, developer, or tester and collaborates via chain-of-thought prompting. The set of LLM agents emulate different software process models and are evaluated against four benchmarks. The impact of individual development activities on code quality is also investigated, revealing the critical role of testing and the beneficial effects of design and code review
Similarly, testing is an important area for applying improved generation methods from models supported by AI. The test case generation and prioritisation presented in [107] addresses the challenges of resource constraints, high costs and inefficiencies in traditional test development methods.
Case studies on AI-augmented MDSE have also been published. [2] explores the use of LLMs to aid in the complex process of re-engineering existing software variants into a software product line (SPL) by identifying features to represent variability and commonalities. It describes the methodology used to prompt the LLMs, the results obtained, and the limitations encountered. There is a potential in assisting domain analysts and developers in transforming software variants into SPLs, but LLMs are limited in managing large variants or complex structures. [40] presents a novel software architecture designed to address the increasing complexity of modern system development within a DevOps framework. The architecture integrates MDSE and AI techniques to enhance the entire software lifecycle, from requirements engineering to monitoring and maintenance. The architecture is evaluated through multiple industrial case studies.
@@
In general, the potential of using AI for MDSE can be summarised as follows: MDSE provides the formal structure and ’guardrails’ that are required for AI to operate effectively and reliably. In turn, AI provides the intelligence and flexibility needed to automate the most challenging and creative stages of the MDSE lifecycle. To realise these prospects in industrial SE practices, the following opportunities require broader uptake: